# DriveSOTIF: Advancing Perception SOTIF Through Multimodal Large Language Models

### v1

@@ IV-C Dataset Overview
Number of questions 4455 1115
Number of answers 44550 11150
Number of captions 891 223
Average question length (words) 5.30 5.37
Average answer length (words) 3.51 3.44
@@
As shown in Table II, our generated dataset is highly related to perception-related SOTIF problem and and can be utilized to enhance MLLMs for VQA/image captioning tasks in AD. Furthermore, our dataset features an average caption length of 18.22 to 19.22 words, nearly double the average length of 10.5 words per caption in the COCO Caption [26] dataset. This increased length provides additional context and could potentially enhance model performance.
IV-D Evaluation Metrics
For evaluation, we primarily focus on two key aspects: the similarity between model-generated answers and human-generated answers, and the compatibility between the image and the model-generated response.
In image captioning, we used standard metrics such as BLEU, METEOR, ROUGE, CIDEr [30], and SPICE [31]. CIDEr calculates the average cosine similarity between candidate and reference sentences, while SPICE evaluates precision and recall based on scene graphs. Together, these metrics assess the accuracy, fluency, and relevance of the generated captions, providing a comprehensive evaluation of the model’s ability to describe visual scenes. For the VQA task, we relied on BLEU and ROUGE scores to measure how accurately and contextually the models answered visual questions. These metrics highlight the models’ proficiency in generating meaningful and relevant responses based on the given visual inputs.
V Model Benchmarking and Fine-tuning
V-A Benchmarking
For image captioning benchmarking, we evaluate the Blip [32] and Blip2 models [33] (pre-trained on the COCO Caption dataset [26]) across various model sizes. For VQA tasks, we benchmark a total of five open-sourced models, including LLaVA 1.5 [34], LLava-next[35], LLaVA-Video[36], Idelics2 [37] and newly released Qwen2-VL [38] that is developed based on Qwen2 [39].
These models are classified by architecture, model size, and quantization method, focusing on three specific variants: 2B, 7/8B, and 13/34B. The 2B variant is particularly suitable for mobile device deployment due to its lightweight and performance. The 7/8B model is well-suited for on-vehicle computing, where there is a need for both performance and efficiency. The 13/34B variant, due to its larger size and resource requirements, is more appropriate for cloud-based deployments. The model benchmarking and fine-tuning presented are completed on a desktop equipped with an Intel i9-10900K, 128GB of memory, and two Nvidia RTX 3090 GPUs.
V-B Model Fine-tuning
@@
For the image captioning task, we fine-tune the Blip [32] and Blip2 [33] models using the LAVIS framework [42]. In the VQA tasks, we leveage LLaVA [34] and LLaMA-Factory [43] frameworks to fine-tune LLaVA 1.5 and Qwen2-VL family models due to their high performance scores when compared with other models.
To ensure the reproducibility of our results, we maintain consistent settings across fine-tuning and evaluation phases. As shown in Table IV, we use identical random seeds and hyperparameters, such as learning rate (LR), temperature, for all models.
TABLE IV: Settings for Fine-tuning and Evaluation
Parameter Value Parameter Value
Image Captioning
Max Epoch 5 LR Scheduler Linear Warmup Cos
Max Length 50 Initial LR 1e-5
Min Length 8 Minimum LR 0
Number of Beams 5 Warmup LR 1e-8
Image Size 364 Warmup Steps 1000
Freeze ViT True Weight Decay 0.05
VQA
Learning Rate 5e-5 Max New Tokens 512
Cutoff Length 1024 Temperature 0.95
Top-p 0.7 Quantization BitsandBytes
V-C Results and Analysis
V-C1 Qualitative Analysis
Qualitatively, we illustrate the results using three representative images from the test split, as depicted in Table V, where . These results include the labeled ground truth, the predictions generated by the baseline models, and those produced by the fine-tuned models. For image captioning, baseline model is BLip2 6.7B while fine-tuned Blip2 6.7B . For VQA, responses from raw and fine-tuned QWEN2VL-7B are presented.
From Figure V, it can be seen that baseline models for image captioning tend to give short, generic descriptions like “a car driving on a wet road” or “a car driving on a foggy road.” These captions, while accurate, often miss important details that add depth to the scene, such as specific road conditions or potential hazards. On the other hand, fine-tuned models offer much more detailed and context-aware descriptions. For example, instead of just noting the basic scene elements, they might say, “a driver’s view through a windshield on a rainy day, with red traffic lights ahead,” drawing attention to factors like visibility and potential obstructions. In foggy conditions, they might highlight “low visibility”, offering a clearer understanding of the challenges faced by the driver.
In the VQA tasks, the baseline models generally perform adequately when answering simple, close-ended questions but tend to struggle with more complex, open-ended queries. In contrast, the fine-tuned models demonstrate a significant improvement in handling open-ended questions by providing more detailed and context-aware answers.
For instance, in response to a question about rain’s impact on perception, the fine-tuned model identifies challenges like “reflections and distortions caused by the rain on the windshield,” offering a deeper understanding of the driving environment.
In more complex scenarios, such as an overturned vehicle, the fine-tuned model provides more thorough answers, highlighting issues like blocked lines of sight and shadow zones that could complicate the vehicle’s response. These detailed insights showcase the fine-tuned model’s superior ability to handle complex, safety-critical situations.
@@ V-C1 Qualitative Analysis
Task Type Image 1 Image 2 Image 3
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
@@
IC Base: a view of a car driving on a wet road a car driving on a foggy road a car that is upside down on the side of the road
@@
FT: a driver’s view through a windshield on a rainy day, with red traffic lights visible on the road ahead a driver’s view through a car’s windshield on a foggy road, with low visibility and limited visibility a car flips over on a busy road, causing traffic congestion and hindering visibility for drivers on a busy city street
@@
GT: View from a vehicle’s dashboard camera showing a rainy and blurred roadway scene with visible windscreen wiper, a truck ahead, and blurry tail lights Intense fog conditions on a road pose a challenge for drivers, with only faint outlines of vehicles ahead visible An overturned car in the middle of a busy street highlights potential hazards in urban traffic conditions
VQA Q1: Is there precipitation affecting visibility in the scenario depicted? Is the visibility reduced in this driving scenario due to fog or mist? Is the overturned vehicle a key object in the scene?
Base: Yes Yes Yes
FT: Yes Yes Yes
GT: Yes Yes Yes
@@
Q2: What weather condition is depicted in the image that could affect perception systems in vehicles? What environmental factor is degrading the perception ability in this image? What perception-related SOTIF risk is present due to the vehicle’s unusual position in the scene?
@@
Base: rainy weather, which can affect the perception systems by reducing visibility, making it more difficult for the systems to accurately detect and track objects. The image shows that the environmental factor degrading perception ability is fog camera’s field of view may not capture the full extent of the scene
@@
FT: The image shows heavy rain conditions, which can significantly degrade the performance of perception systems in vehicles by scattering and reflecting light The perception ability is degraded by heavy fog, which significantly reduces visibility on the road. The overturned vehicle poses a risk as it is not commonly expected to be in that position on the road, which can lead to misclassification or delayed recognition by autonomous driving systems.
@@
GT: The image shows heavy rain conditions which can degrade the visibility and accuracy of vehicle perception systems by scattering light and obscuring road markings and other vehicles. The environmental factor shown in the image is fog, which significantly reduces visibility and therefore degrades the perception ability of drivers and automotive sensors. The overturned vehicle presents a high SOTIF risk by unexpectedly altering the driving environment, which could lead to inappropriate responses from an autonomous driving system unprepared for such irregular situations.
Q3: What specific aspect of the environment in this image can autonomously driven vehicles struggle to interpret correctly? What manual safety precaution should be taken by drivers in foggy conditions like in the image? How could the presence of other vehicles near the overturned car influence the autonomous vehicle’s response?
Base: The image shows a road with rain, which can make it difficult for autonomous vehicles to interpret the environment correctly. Rain can cause the road surface to become slippery, making it more difficult for vehicles to maintain traction and control. Additionally, rain can reduce visibility, making it harder for vehicles to see obstacles and other vehicles on the road. Decrease speed to allow for better reaction time. Increase following distance to give enough time to react if the vehicle in front stops suddenly. Use headlights to improve visibility, but avoid high beams as they can blind other drivers. Use fog lights if available for additional illumination. Keep windshield wipers off to avoid blurring visibility. Ensure mirrors are clear and properly adjusted. Avoid sudden braking, acceleration, or turning. Use turn signals to inform other drivers of your intentions. Stay in your current lane and avoid lane changes unless absolutely necessary. The presence of other vehicles near the overturned car could influence the autonomous vehicle’s response in several ways. First, the autonomous vehicle may need to adjust its speed and direction to avoid colliding with the overturned car and other vehicles on the road. Second, the autonomous vehicle may need to communicate with other vehicles on the road to ensure that they are aware of the situation and take appropriate actions. Finally, the autonomous vehicle may need to coordinate with emergency services to ensure that they are aware of the situation and can respond accordingly.
FT: The reflections and distortions caused by the rain on the windshield can lead to misinterpretation of the road condition and the position of other vehicles or pedestrians, which are critical for safe navigation Drivers should reduce their speed, increase following distances, use fog lights if available, and ensure their vehicle’s headlights and windshield are clean for better visibility. Other vehicles could block line-of-sight or create shadow zones, complicating the vehicle’s ability to assess the situation and determining the appropriate response.
GT: Autonomously driven vehicles can struggle to correctly interpret lane markings and other road signs obscured by water and reflections on the wet road surface, increasing the risk of navigation errors. Drivers should reduce their speed, use fog lights if available, maintain a safe following distance, and stay focused on the road to adjust quickly to potentially unseen hazards. Other vehicles close to the overturned car might limit maneuvering options for an autonomous vehicle and complicate the scene understanding, necessitating advanced decision-making algorithms that incorporate predictions of other drivers’ behaviors.
V-C2 Quantitative Analysis
The quantitative results, shown in Tables V-C2 and V-C2, demonstrate that fine-tuned models consistently outperform their baseline counterparts across all evaluation metrics.
@@ V-C2 Quantitative Analysis
Method Size Trainable Parameters (M) BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr SPICE
Baselines
Blip [32] Small 223.97 0.348 0.219 0.132 0.082 0.134 0.315 0.138 0.086
Large 446.29 0.349 0.227 0.139 0.083 0.137 0.324 0.140 0.091
@@
6.7B 107 0.398 0.270 0.169 0.106 0.146 0.338 0.176 0.096
Fine-Tuned Models
Blip [32] Small 223.97 0.659 0.464 0.318 0.217 0.208 0.414 0.411 0.143
Large 446.29 0.671 0.463 0.312 0.206 0.210 0.399 0.410 0.145
@@
6.7B 107 0.723 0.509 0.353 0.248 0.225 0.421 0.536 0.155
TABLE VII: VQA Scores Comparison: Baseline vs. Fine-tuned Models
Method Size Quan. BLEU-4 ROUGE-1 ROUGE-2 ROUGE-L
Base FT Base FT Base FT Base FT
LLaVA 1.5 [34] 7B - 18.95 46.06 26.09 40.40 5.70 13.14 14.80 30.86
13B 8 bit 18.83 46.50 26.21 41.27 5.69 14.47 14.96 31.85
@@
7B - 25.55 46.36 29.38 40.63 7.14 13.27 18.50 30.80
In image captioning, the fine-tuned Blip2 6.7B model obtained the best results, achieving ROUGE-L, CiDEr, and SPICE scores of 0.421, 0.536, and 0.155, respectively. This marks an increase of 24.56%, 204.54%, and 61.46% compared to the baseline version. Similar improvements were observed in METEOR, ROUGE, CIDEr, and SPICE scores, indicating the models’ enhanced ability to generate more accurate and contextually relevant descriptions of complex scenes. Notably, the fine-tuned Blip2 2.7B model achieved comparable performance, highlighting the importance of balancing model size with computational efficiency and performance needs.
For the VQA task, among baseline models, the newly released Qwen2-7B achieved the highest scores in BLEU, and ROUGE metrics. Additionally, it can be seen that larger models generally achieve better performance. For example, the 34B version of LLaVA-Next, implemented using 4-bit precision, outperformed its 13B and 7B variants. A similar trend is observed when comparing the Qwen2-VL and LLaVA 1.5 models. For the fine-tuned models, the LLaVA 1.5 - 13B model deployed with 8-bit precision achieved the highest scores across all metrics, which is expected given its larger size. Interestingly, the Qwen2-VL 2B model delivered similar level of performance with only 1/7 the number of parameters, making it a highly efficient option for deployment.
Similar to image captioning, fine-tuning also led to significant performance gains in VQA tasks. The BLEU-4 score for the LLaVA 1.5 model 13B version increased by 146.95% from 18.83 to 46.50, significantly improving the model’s accuracy and contextual understanding in handling open-ended questions. Similar improvements can be seen across all other metrics, showing that the fine-tuned models consistently generate more relevant and contextually appropriate responses.
V-C3 Discussion
Based on the results from both the quantitative and qualitative analyses, it is evident that fine-tuning leads to significant improvements in model performance across image captioning and VQA tasks. Fine-tuned models consistently generate more accurate, detailed, and contextually relevant outputs compared to their baseline counterparts.
Additionally, our findings demonstrate that the DriveSOTIF dataset, despite its relatively small size of 1,114 images, can effectively enhance the performance of MLLMs through SFT. This highlights the potential of using focused, domain-specific datasets to significantly improve model performance. Lastly, the proposed dataset generation approach can be adopted in other autonomous driving tasks, offering an efficient pathway for adopting MLLMs into specialized applications.
VI Experiments
VI-A Model Deployment and Testing
As illustrated in Figure 3, common deployment framework for MLLM models include Huggingface, vLLM [44], and TensorRT-based methods [45]. These models can be deployed locally on devices, in the cloud, or through multi-access edge computing (MEC) or edge computing setups. At a high level, the deployment process can be simplified to an API request/response service, as shown in Figure 3. In this setup, input images and text are sent to the API, where deployed MLLM processes them and returns the response.
In this research, the fine-tuned versions of Qwen2-VL 2B and 7B models were chosen for deployment due to their efficient performance and compact size. These models were deployed using the Lamma-Factory framework [43] through Huggingface, with a batch size set to 1. Both models are merged with their corresponding fine-tuned adaptive weights and deployed on a single RTX 4090 GPU.
For testing, 111 images from the test split of DriveSOTIF dataset were used. Each image was paired with the following input prompt: “We are studying perception-related SOTIF problems for autonomous driving vehicles. What are the SOTIF risks of this image, and what are the reasons and suggested actions? Provide an informative, natural, human-like, and accurate response. Do not make up information.”
Figure 3: Overview of multi-model LLM deployment.
Model Mem. [GB] Quant. Mean [s] Median [s] Std Dev [s] Min [s] Max [s] 2B 8.0 - 0.82 0.80 0.11 0.59 1.23 6.3 8bit 6.45 6.22 0.85 4.68 8.80 4.8 4bit 2.21 2.17 0.36 1.44 3.30 7B 19.5 - 2.78 2.33 1.18 0.98 6.02 13.2 8bit 10.04 8.44 4.31 5.00 23.42 9.2 4bit 3.57 3.09 1.55 1.44 7.98
TABLE VIII: Processing time statistics for 2B and 7B models.
Figure 4: Processing time per image
The peak GPU memory usage and processing time per image are recorded and plotted in Figure 4.
The statistical analysis is performed, and the result is presented in Table VIII. Based on our results, quantization primarily reduces GPU memory usage but does not improve real-time performance. This is reasonable since only the pre-trained LLM is quantized, while other computations still operate with higher precision. As a result, for the Qwen2-VL 2B model, GPU memory usage decreases from 8.07 GB (full precision) to 6.33 GB (8-bit) and further down to 4.78 GB (4-bit). A similar trend can be observed in the 7B model.
In terms of real-time performance, for the Qwen2-VL 2B model, the non-quantized version had an average processing time of 0.82 seconds per image, with a median of 0.80 seconds and a standard deviation of 0.11 seconds. When the model was quantized to 8-bit precision, the processing time increased to 6.45 seconds per image. In contrast, using 4-bit precision reduces it to 2.21s per image. A similar pattern can be seen in the Qwen2-VL 7B variant.
The significant slowdown for 8-bit occurs because the consumer-grade GPU we used isn’t optimized to handle calculations with lower-bit-width tensors directly. As a result, the system has to perform extra steps to convert these lower-bit tensors back into float values before performing the necessary computations [46]. These additional steps introduce significant overhead, dramatically slowing down the process. In contrast, 4-bit quantization is slightly faster than 8-bit. This happens as 8-bit often requires intermediate conversions, while 4-bit tensors can be packed more efficiently, leading to reduced memory bandwidth overhead and improved performance. Nonetheless, with dedicated hardware and further optimization and support of the deployment pipeline, this slowdown could be mitigated.
VI-B Real World Data Evaluation

### v2

@@ III-C Dataset Overview
Number of questions 5,025 555
Number of answers 50,250 5,550
Number of captions 1,005 111
Average question length (words) 11.9 11.8
Average answer length (words) 20.5 21.0
@@
Dataset Content Distributions
Difficulty Distribution
 Easy questions 2,224 (39.9%)
 Medium questions 2,243 (40.2%)
 Hard questions 1,113 (19.9%)
Question Type Distribution
 Close-ended questions 2,549 (45.7%)
 Open-ended questions 3,031 (54.3%)
Expected Answer Type Distribution
 Analysis 2,014 (36.1%)
 Yes/No & multiple choice 1,568 (28.1%)
 Recommendation 995 (17.8%)
 Count 947 (17.0%)
 Identification 56 (1.0%)
Generation Information
Average Generation Attempts
 Caption generation 1.16 attempts/item
 Question generation 1.23 attempts/item
 Answer generation 1.72 attempts/item
Success Rates by Attempt Number
 Caption generation 90.1% / 8.3% / 1.6%
 Question generation 82.7% / 13.5% / 3.9%
 Answer generation 71.0% / 21.4% / 7.6%
III-D Evaluation Metrics
To comprehensively evaluate model performance, we focus on two primary aspects: the similarity between model-generated answers and human-annotated ground truth, and the compatibility between generated responses and visual content.
For image captioning, we adopt established metrics including BLEU, METEOR, ROUGE, CIDEr [29], and SPICE [30]. CIDEr computes the average cosine similarity between candidate and reference captions, while SPICE measures precision and recall based on scene graphs. Together, these metrics provide a holistic assessment of caption accuracy, fluency, and contextual relevance.
For close-ended VQA tasks, we use accuracy scores, which measure the percentage of correct answers generated by the model. For open-ended VQA, where no ground truth exists, we follow a rubric-based approach, leveraging an LLM-as-judge [31] to score answers. Specifically, we use GPT-4.1 to evaluate responses based on the question, ground-truth answer, input image, and the model’s output. Each response is scored from 1 to 5 across multiple criteria: relevance, trustworthiness, clarity, and coherence[32]. The evaluation is repeated three times with a temperature of 0.7 to maximize consistency with human judgment.
This evaluation strategy enables nuanced assessment of both factual correctness and the quality of higher-level reasoning in perception-related SOTIF scenarios.
IV Model Benchmarking and Fine-tuning
IV-A Benchmarking
For image captioning benchmarking, we evaluate the Blip [33] and Blip2 models [34] (pre-trained on the COCO Caption dataset [27]) across various model sizes. For VQA tasks, we benchmark a total of Four open-sourced models, including LLaVA 1.5 [35], Qwen2/2.5-VL [36, 37], and InterVL3 [38].
These models are classified by architecture, model size, and quantization method, focusing on two specific variants: 2/3B and 7/8B. The 2/3B variant is particularly suitable for mobile device deployment due to its lightweight design and efficient performance. The 7/8B model is well-suited for on-vehicle computing, where high computational capability is required. By evaluating across this spectrum, we aim to understand trade-offs between accuracy, resource demands, and deployment feasibility for perception-related SOTIF applications in autonomous driving.
IV-B Model Fine-tuning
@@
For image captioning, we fine-tuned Blip and Blip2 models using the LAVIS framework [41]. For VQA, we utilized LLaVA [35] and LLaMA-Factory [42] frameworks to fine-tune LLaVA 1.5, Qwen2/2.5-VL, and other leading vision-language models. Hyperparameters, random seeds, and training configurations were kept consistent across experiments for fair comparison (details in S.II of the supplemental material).
@@
Task Type Image 1 Image 2
[Uncaptioned image] [Uncaptioned image]
@@
IC Base: a view of a car driving on a wet road a car that is upside down on the side of the road
@@
FT: a driver’s view through a windshield on a rainy day, with red taillights visible ahead a car flips over on a busy road, causing traffic congestion and hindering visibility for drivers on a busy city street
@@
GT: View from a vehicle’s dashboard camera showing a rainy and blurred roadway scene with visible windscreen wiper, a truck ahead, and blurry tail lights An overturned car in the middle of a busy street highlights potential hazards in urban traffic conditions
@@
VQA Q: What weather condition is depicted in the image that could affect perception systems in vehicles? What perception-related SOTIF risk is present due to the vehicle’s unusual position in the scene?
@@
Base: rainy weather, which can affect the perception systems by reducing visibility, making it more difficult for the systems to accurately detect and track objects. camera’s field of view may not capture the full extent of the scene
@@
FT: The image shows heavy rain conditions, which can significantly degrade the performance of perception systems in vehicles by scattering and reflecting light The overturned vehicle poses a risk as it is not commonly expected to be in that position on the road, which can lead to misclassification or delayed recognition by autonomous driving systems.
@@
GT: The image shows heavy rain conditions which can degrade the visibility and accuracy of vehicle perception systems by scattering light and obscuring road markings and other vehicles. The overturned vehicle presents a high SOTIF risk by unexpectedly altering the driving environment, which could lead to inappropriate responses from an autonomous driving system unprepared for such irregular situations.
@@
Method Size BLEU-1 BLEU-2 BLEU-3 BLEU-4 METEOR ROUGE-L CIDEr SPICE
Base FT Base FT Base FT Base FT Base FT Base FT Base FT Base FT
Blip Small 0.348 0.659 0.219 0.464 0.132 0.318 0.082 0.217 0.134 0.208 0.315 0.414 0.138 0.411 0.086 0.143
Large 0.349 0.671 0.227 0.463 0.139 0.312 0.083 0.206 0.137 0.210 0.324 0.399 0.140 0.410 0.091 0.145
@@
6.7B 0.398 0.723 0.270 0.509 0.169 0.353 0.106 0.248 0.146 0.225 0.338 0.421 0.176 0.536 0.096 0.155
TABLE VII: VQA Scores Comparison: Baseline vs. Fine-tuned Models
Model Size Accuracy (%) Relevance Trustworthiness Clarity Coherence Overall
Base FT Base FT Base FT Base FT Base FT Base FT
LLaVA 1.5 7B 57.51 64.06 3.88 4.01 4.04 4.44 4.69 4.85 4.68 4.89 4.32 4.54
@@
7B 56.89 62.13 4.24 4.14 4.32 4.51 4.81 4.90 4.79 4.90 4.54 4.61
@@
2B 58.01 64.84 4.22 3.92 4.21 4.34 4.74 4.80 4.76 4.82 4.48 4.47
8B 59.12 64.49 4.56 4.12 4.70 4.53 4.94 4.90 4.96 4.92 4.79 4.62
IV-C Results and Analysis
IV-C1 Qualitative Analysis
Table V presents representative examples illustrating how fine-tuning improves model performance on both image captioning and VQA tasks. Baseline captioning models tend to generate short, generic descriptions (e.g., “a car driving on a wet road”) that miss key scene details. After fine-tuning, models produce richer, context-aware captions such as “a driver’s view through a windshield on a rainy day, with red taillights visible ahead,” capturing both environmental context and potential hazards.
For VQA, baseline models typically succeed on simple, factual questions but often fail to provide detailed, situation-specific reasoning. Fine-tuned models demonstrate better understanding of complex queries. For example, when prompted about rain-related perception issues, the fine-tuned model highlights specific challenges like reflections and distortions, offering a deeper analysis of SOTIF risks. In unusual scenarios—such as an overturned vehicle—fine-tuned models identify relevant perception challenges (e.g., blocked sightlines or misclassification risks) that baseline models may overlook.
IV-C2 Quantitative Analysis
Fine-tuned models show consistent and significant improvements across standard metrics. In image captioning, the fine-tuned Blip2 6.7B model achieves ROUGE-L, CIDEr, and SPICE scores of 0.421, 0.536, and 0.155, representing increases of 24.6%, 204.5%, and 61.5% over the baseline, respectively. Similar improvements are observed across BLEU, METEOR, and other benchmarks, indicating enhanced contextual relevance and descriptive accuracy. Notably, the smaller Blip2 2.7B model also performs competitively, highlighting the value of balancing model size and efficiency.
For VQA, larger models generally outperform smaller ones, with the 8B variant of InternVL3 reaching the highest baseline accuracy at 59.1%. Fine-tuning yields substantial accuracy gains for all models: for instance, InternVL3 2B rises from 58.0% to 64.8%. Trustworthiness, clarity, and coherence metrics—scored by LLM-as-judge—also improve, often by more than 0.6 points on a 5-point scale. Of note, fine-tuned Qwen2-VL 2B achieves competitive performance (61.01% accuracy) with only a fraction of the parameters compared to larger models. While accuracy and most quality metrics improve substantially, some metrics such as “relevance” show smaller gains, that even baseline models can identify broad risk factors. However, they lack the detailed, domain-specific knowledge required for precise safety recommendations.
Both qualitative and quantitative results confirm that supervised fine-tuning on the DriveSOTIF dataset significantly enhances model performance for perception-related SOTIF scenarios. The domain-specific dataset enables models to better interpret complex, safety-critical driving situations, supporting more robust and reliable reasoning for autonomous driving applications.
V Experiments
V-A Model Deployment and Real-time Performance Evaluation
MLLMs can be deployed across various environments, including local devices, cloud platforms, and edge computing systems such as multi-access edge computing (MEC). Common deployment frameworks include Huggingface, TensorRT-based solutions [43], and the vLLM framework [44], which was selected for this research due to its efficiency and support for advanced scheduling strategies.
The deployment setup follows a simplified API request-response architecture: input images and text are sent to the deployed model via API calls, and the model returns the corresponding output. For this work, compact, fine-tuned MLLMs were used to ensure fast and resource-efficient inference. These models were merged with their respective adaptive weights and deployed on two platforms: NVIDIA Jetson Orin and a consumer-grade RTX 3090 GPU.
Two inference strategies were considered. Sequential inference processes inputs one at a time as they arrive. This provides deterministic, per-sample latency—an essential requirement for real-time, safety-critical applications like autonomous driving. Continuous batching, by contrast, dynamically aggregates requests and schedules token-level processing across partially completed sequences. Instead of traditional static batching, vLLM uses iteration-level scheduling and PagedAttention to refill GPU slots as soon as they are available. This maximizes GPU utilization and minimizes token-level idle time, resulting in higher throughput without sacrificing latency [44].
V-A1 Sequential Processing
We tested this mode using 111 images from DriveSOTIF, each paired with a prompt for SOTIF risk analysis. Generation time for each image were recorded and aggregated in Table V-A1. GPU VRAM usage was measured via HuggingFace APIs due to vLLM’s memory pre-allocation behavior.
[b]s
TABLE VIII: Real-time performance across hardware platforms and quantization levels.
RTX 3090 Jetson AGX Orin1
Model Size Quant RAM Usage Mean Median Std Mean Median Std
[GB] [s] [s] [s] [s] [s] [s]
Qwen2-VL 2B - 4.74 0.82 0.80 0.11 3.70 3.60 0.27
FP8 2.91 0.74 0.73 0.03 2.93 2.89 0.18
4bit 2.05 0.87 0.86 0.04 3.56 3.49 0.23
@@ V-A1 Sequential Processing
2B - 4.14 1.09 1.07 0.16 4.21 4.14 0.40
  • 1
    We utilized vLLM docker for this testing. Orin is set to MAXN. We encountered system throttling warnings due to over-current.
  • 2
    LMDeploy is used for InternVL3 deployment.
The peak GPU memory usage and processing time per image are recorded. The statistical analysis is performed, and the result is presented in Table V-A1. Based on our results, quantization reduces GPU memory usage while improving real-time performance. Specifically, Qwen2-VL (2B) in 4-bit quantization achieves the lowest memory footprint (2.05 GB) but with a slight latency increase compared to FP8. FP8 quantization yields the best speed on Jetson Orin, reducing inference time by approximately 21% relative to the full-precision model. For larger models like Qwen2.5-VL (3B), 4-bit quantization significantly reduces memory usage from 7.64 GB to 2.90 GB and improves inference speed on both hardware platforms, achieving over 1.3× speedup on Jetson Orin.
Comparing across models, InternVL3-1B consistently exhibits strong performance with the lowest latency and compact memory usage, making it suitable for deployment in resource-constrained environments. However, as the model size increases (e.g., InternVL3-2B), the latency grows accordingly, especially on edge devices.
V-A2 Continuous Processing for high-speed driving
To evaluate continuous inference, we replayed four 30 Hz video streams of high-speed driving scenarios, triggering model inference on every frame. Each frame generates a request to the InternVL3-1B vision-language model, which is deployed on an RTX 3090 with continuous batching enabled  [^4] Please refer to Section V of the supplemental material for results on RTX4090 and duel RTX4090 setup. The system supports multiple requests in parallel, and both queuing and auto-timeout mechanisms are in place: when more requests arrive than the hardware can process immediately, excess requests are placed in a first-in, first-out (FIFO) queue, and any request waiting or processing beyond a set timeout threshold is automatically dropped. We varied the maximum number of in-flight (active) requests—from 1 up to 30 (concurrency)—and, for each setting, recorded the API response times and the number of active requests throughout the experiment (see Figure 3).
At low concurrency levels (1, 3, 5), response times remained stable and low, averaging around 0.5 to 0.7 seconds per frame. As concurrency increased past 10, response times began to rise rapidly, reaching an average of 3.5 seconds per frame at the highest level tested (30). This sharp jump signals that the model and GPU are saturated, leading to longer waits and increased resource contention.
We visualized these dynamics by jointly plotting queue wait times and API response times for each concurrency setting (Figure 4). At low concurrency, the system processes requests quickly once they reach the model, but incoming requests rapidly accumulate in the queue, leading to prolonged wait times and, frequently, discarded requests due to timeouts. As concurrency increases, response times become more variable and overall latency rises, but queue wait times decrease since more requests are processed in parallel. In contrast, with very low concurrency,
This illustrates a fundamental tradeoff between response time and queue waiting time with various concurrency levels. High concurrency leads to longer per-request processing times but keeps queue wait times low. Low concurrency minimizes processing latency for individual requests but causes significant queuing and dropped requests under heavy load. For real-time applications such as autonomous driving, it is essential to maintain concurrency at a moderate level to balance response latency and throughput, ensuring timely and reliable perception without excessive delays or data loss.
Figure 3: Average generation time per image in continuous inference mode.
Figure 4: Tradeoff between API reponse time and queue waiting time in continuous inference mode.
V-B Real World Data Evaluation